[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Nov 07, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Mon Nov 7 15:02:23 UTC 2016


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. TAFB forecasters had difficulties accessing the Grunt workstation so  
they called the SDM. An NCO SysAdmin diagnosed the problem as a local  
hardware failure. Reseating the USB cables in the back of the machine  
restored grunt to full functionality.

2. N-AWIPS issues: Error messages "failure to send to FTP Gate" continue to  
appear in all N-AWIPS launch scripts for all TAFB products. Matt Sardi will  
look into turning off this old, secondary connection in all send scripts.
HSU:
1. The times in the TWO (tropical weather outlook) products had to be  
entered by hand due to a known bug in the perl script when the time "falls  
back" from Daylight Savings Time to EST.
Other items:

NHC RFCs:
1. RFC #208 - ldm_ensure v1.1.0 - Copy the 'ldm_ensure' script from  
dbn2.nhc.noaa.gov to logger.nhc.noaa.gov. This script has 2 primary checks  
for the ldm. Only the first check (that the ldm is running and processing  
files) applies to logger, so the second (specific to NOAAports or upstream  
other upstream ldms) will be commented out. To be implemented on Monday  
November 7 at 1200Z (7 am EST).

2. RFC #209 gempak/NAWIPS v7.2.0 - Update $NCSITE/datatype.tbl and  
miscset.tbl tables to display ASCATD High Resolution data on operational N-  
AWIPS systems. Data are currently being ingested via DBNet. To be  
implemented on Tuesday, November 8 at 1600Z (11 am EST).

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161107/2c8d2ef6/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list